On 1 August 2016, NBS Ltd was incorporated and a prospectus was issued inviting applications for 200 000 ordinary shares to the public at an issue price of $8, payable as follows:

$4 on application (due by closing date of 1 October)

$2 on allotment (due 1 November)

$2 on final call/calls (to be determined by the directors)

By 1 October, applications had been received for 230 000 ordinary shares of which applicants for 50 000 shares forwarded the full $8 per share, applicants for 100 000 shares forwarded $6 per share and the remainder forwarded only the application money.

At a directors’ meeting on 7 October, it was decided to allot shares in full to applicants who had paid the either $8 or $6 on application, to reject applications for 20 000 shares and to proportionally allocate shares to all remaining applicants. According to the company’s constitution, all surplus money from application can be transferred to Allotment and/or Call accounts. All outstanding allotment money was received by the due date. Legal costs of forming the company were $2 000 and were paid on 11 October. Share issue costs of $10 000 were also paid on the same date.

Required:

1. Prepare general journal entries to record the above transactions of NBS Ltd. Narrations are not required.

2. Calculate the amount of share capital in the statement of financial position of NBS Ltd as at 30 June 2017 (Show all workings.)
Answer #1

SOLUTION-

date particulars dr cr
1/10/2016 bank a\c dr. 1320000
To share application 1320000
(application money recd)
7/10/2016 share application a\c dr. 1320000
to bank (20000*4) (for rejection) 80000
To share capital(200000*4) (i) 800000
To share allotment 340000
to share first call 100000

(adjustment of money)

11/10/2016 legal cost a/c dr. 2000
share issue cost a/c dr. 10000
to cash 12000
(expenses)
1/11/2016 share allotment a/c dr (200000*2) 400000
to share capital (200000*2) (ii) 400000
(allotment due)
bank a/c dr 60000
to share allotment 60000
(allotment recd)
before june30,2017 share first call a/c dr. 400000
to share capital (iii) 400000
(first call due)
bank a/c dr. 300000
to share first call 300000
(first call recd.)

working notes

1.application rejection money=20000*4=80000

2. for share allotment in second entry=share allotment money recd. in advance from 50000shares+share allotment money recd. in advance from 100000 shares+adjustment of proportionally allocated shares=50000*2+100000*2+10000*4=340000

3.for allotment money due=200000*2=400000

4.for first call due=200000*2=400000

5.for amount recd. on share allotment=400000-340000(already recd.)=60000

7.for first call recd.=400000-100000(already recd.)=300000

3.in the subsequent entries money recd. in advance has been adjusted.

AMOUNT OF SHARE CAPITAL

=(iI+(ii)+(iii)=800000+400000+400000=1600000

as there is no calls in arrear you can also show it as 200000*8=1600000

it has been assumed that first call is also made before 30th june 2017